# Breaking Imitation Bottlenecks: Reinforced Diffusion Powers Diverse Trajectory Generation

### v2

@@ Abstract
Existing end-to-end autonomous driving (E2E-AD) methods predominantly rely on single expert demonstrations through imitation learning, often leading to conservative and homogeneous driving behaviors that struggle to generalize to complex real-world scenarios. In this work, we propose DIVER, a novel E2E-AD framework that combines diffusion-based multi-mode trajectory generation with reinforcement learning to produce diverse, safe, and goal-directed trajectories. First, the model conditions on map elements and surrounding agents to generate multiple reference trajectories from each ground-truth reference trajectory that overcome the inherent limitations of single-mode imitation. Second, we treat the diffusion process as a stochastic policy and employ Group Relative Policy Optimization (GPRO) objectives to guide the diffusion process. By optimizing trajectory-level rewards for both diversity and safety, GRPO directly mitigates mode collapse and enhances collision avoidance, encouraging exploration beyond expert demonstrations and ensuring physically plausible plans. Furthermore, to address the limitations of L2-based open-loop metrics in capturing trajectory diversity, we propose a novel trajectory diversity metric to evaluate the diversity of multi-mode predictions. Extensive experiments on the closed-loop NAVSIM and Bench2Drive benchmarks, as well as the open-loop nuScenes dataset, demonstrate that DIVER significantly improves trajectory diversity, effectively addressing the mode collapse problem inherent in imitation learning.
@@ 1 Introduction
Figure 1: (a) Imitation-based Single-Mode Trajectory Planning  [1, 2, 3, 4, 5, 6, 7] predicts deterministic trajectories but lacks action diversity, leading to potential safety risks. (b) Imitation-based Multi-Mode Trajectories Planning  [3, 8, 9, 4] fails to address the diversity loss in imitation learning end-to-end autonomous driving, leading to mode collapse. The generated multi-mode trajectories overly depend on a single GT trajectory, ultimately clustering around it. (c) The proposed DIVER framework adopts reinforced diffusion for multi-mode trajectory generation, encouraging the ego-vehicle to produce diverse driving behaviors instead of rigidly following a single expert.
@@
In recent years, autonomous driving has entered a new phase, with significant progress in end-to-end autonomous driving (E2E-AD) [10, 11]. E2E-AD systems integrate traditionally independent tasks—such as 3D object detection, multi-object tracking, online mapping, motion prediction and planning—into a unified framework that directly learns driving policies from raw sensor inputs [1, 2, 3, 4, 5, 6]. Among these tasks, planning serves as a core module that not only directly governs ego-vehicle decision-making and control but also provides implicit guidance to upstream perception tasks [1].
@@
Currently, imitation learning (IL) continues to serve as the predominant methodology for planning task in autonomous driving systems [10]. IL enables E2E-AD methods to mimic human expert driving behaviors through supervised learning, typically by minimizing the distance between predicted trajectories and ground truth (GT) trajectories. In the early stages, E2E-AD methods [1, 2, 5, 6] primarily adopted single-mode trajectory, as shown in Figure 1 (a), which predicts a single trajectory through regression. However, it struggles to handle complex traffic scenarios due to neglecting the uncertainties of the driving environment and interactions with other traffic participants. To address the limitations of single-mode trajectory generation, subsequent studies [8, 3, 4, 9] have proposed multi-mode trajectories prediction frameworks, as illustrated in Figure 1 (b). For instance, VADv2 [8] employs probabilistic modeling techniques to capture potential driving intentions and generate multiple proposal trajectories, ultimately selecting the optimal trajectory through a maximum scoring mechanism. However, these approaches [8, 3, 4, 9] still suffer from mode collapse, where the generated multi-mode trajectories exhibit excessive dependence on GT trajectories. This results in multi-mode trajectories distributions that are overly concentrated around the GT, failing to adequately represent the true diversity of driving behaviors. Mode collapse is a form of diversity loss in multi-mode trajectories prediction [10]. Its root cause is the strong reliance of IL methods on a single expert trajectory. Regardless of open-loop or closed-loop training, and whether in single-mode or multi-mode prediction, most IL methods are trained using a single GT expert trajectory as the target, limiting the learning of diverse driving behaviors.
@@
Some autonomous driving methods [12, 13] leverage diffusion models to generate diverse trajectories. Diffusion models [14], as a powerful generative learning framework, have demonstrated strong capabilities in modeling multi-mode behavior distributions for robotic policy learning [15]. Introducing diffusion models into E2E-AD trajectory planning has emerged as a promising research direction. For instance, DiffusionDrive [4] combines truncated diffusion strategies with efficient cascaded diffusion decoders to mitigate the loss of trajectory diversity during IL training. However, since its optimization process still adheres to the IL framework based on a single expert trajectory, the generated multi-mode trajectories ultimately tend to converge toward a single GT. This limitation prevents the model from fully overcoming the inherent constraints of imitation learning.
Although diffusion models [16] demonstrate strong capabilities in multi-mode trajectories generation, their training still primarily relies on maximum likelihood estimation. While conditional constraints can provide some guidance, the lack of explicit supervision for safety often leads to physically infeasible trajectories in real-world scenarios. Recent works such as DeepSeek-R1 [17] have leveraged reinforcement learning (RL) to enhance reasoning capabilities, while AlphaDrive [18] and RAD [19] have validated the effectiveness of RL in E2E-AD. Unlike conventional supervised learning, RL directly evaluates and optimizes generation results based on designed reward functions without requiring explicit labels. Motivated by this, we propose to integrate reinforcement learning with diffusion models by designing diversity and safety rewards as external optimization signals. This approach effectively overcomes the limitations of traditional supervision (e.g., L2 loss) and guides the diffusion model to generate safe, physically plausible, and diverse future trajectories.
@@
In this work, we propose DIVER, a novel multi-mode E2E-AD framework that leverages reinforcement learning to guide diffusion models for generating diverse, feasible trajectories (Figure 1(c)). By leveraging diffusion models, our approach generates multiple plausible future trajectories, capturing a spectrum of driving behaviors while maintaining safety and feasibility. Specifically, we introduce the Policy-Aware Diffusion Generator (PADG), which incorporates map elements and surrounding agents as conditions and generates diverse multi-mode trajectories and multiple reference GTs, guided by a single GT trajectory and trajectory anchors, effectively capturing various driving styles. To ensure both safety and diversity in trajectory planning, we incorporate reinforcement learning to supervise the diffusion process via reward signals, enabling optimization of generated trajectories under diversity and safety constraints. Additionally, to address the misalignment between conventional open-loop metrics (primarily L1Loss) and the intrinsic requirement for diverse trajectory generation, we introduce a novel Diversity Metric to better evaluate the diversity of multi-mode trajectories predictions. Extensive experiments on closed-loop Bench2Drive and NAVSIM, as well as open-loop nuScenes, demonstrate that our DIVER significantly enhances trajectory diversity and planning robustness, effectively addressing the mode collapse issue in imitation learning.
@@
  • We propose a novel Diversity Metric to evaluate multi-mode trajectories generation, providing a more principled way to assess the diversity and effectiveness of generated trajectories compared to existing metrics.
@@ 2.1 End-to-end Autonomous Driving
End-to-end autonomous driving is a fully differentiable machine learning system that takes raw sensor input data and other metadata as prior information, directly outputting control signals or trajectory planning for vehicles [11, 10]. Early works  [20, 21] excluded intermediate tasks, such as perception and motion prediction, which were characterized by limited interpretability and difficulties in optimization. UniAD[1], as a pioneering work, integrates perception tasks and planning tasks via Transformer[22], explicitly supervising 3D object detection [23, 24, 25], multi-object tracking[26], online mapping[27], motion prediction[28], occupancy prediction[29], and planning [30] tasks within a unified model. Subsequently, VAD[2] further explores the application of vectorized scene representation from MapTR [27] to enhance the efficiency of E2E-AD. DualAD [31] disentangles dynamic agents and static elements, compensating for motion effects and flexibly propagating the belief state over time. A series of E2E-AD works [1, 2, 5, 20, 31, 32], represented by UniAD[1] and VAD[2], have adopted single-mode trajectory for planning performance. However, single-mode trajectory planning methods have issues with action diversity and associated safety risks. To address the above limitations, VADv2 [8] has recently shifted to a multi-mode planning framework by scoring and sampling from an extensive fixed vocabulary of anchor trajectories. Based on the design of multi-mode planning, SparseDrive [3] proposes a hierarchical planning selection strategy to select a rational and safe trajectory as the final planning output. Hydra-MDP [9] proposes a universal framework of end-to-end multi-mode planning via multi-target hydra-distillation. By leveraging historical trajectory momentum, MomAD [33] addresses the issue of inconsistent multi-mode trajectories over time. DiffusionDrive [4] leverages diffusion models to capture multi-mode action distributions in an E2E-AD framework. Existing multi-mode trajectories planning methods [8, 3, 9, 33, 4] do not fundamentally resolve mode collapse in imitation learning.
@@ 2.2 Diffusion Models for Autonomous Driving
Diffusion models approximate data distributions through iterative denoising and have demonstrated remarkable performance in image generation tasks [16], showing great potential for behavior modeling and trajectory generation in autonomous driving[34]. DiffScene[35] utilizes diffusion models to synthesize safe scenarios for evaluating autonomous driving safety. DiffBEV [36] adopts conditional diffusion to produce fine-grained BEV representations. MotionDiffuser [37] employs diffusion-based representations for multi-agent trajectory prediction. DiffusionDrive [4] models multi-mode trajectories distributions via a truncated diffusion process. VDT-Auto [38] integrates diffusion Transformers with vision-language models (VLMs) for action generation. MomAD[33] incorporates noise injection and denoising within the perception module to enhance robustness and generalization. DiffAD [39] formulates E2E-AD as a conditional image generation task by unifying heterogeneous driving objectives into a rasterized representation under a diffusion framework. DiFSD [40] introduces both position-level motion diffusion and trajectory-level planning denoising to model uncertainty, thereby enhancing the stability and convergence of the entire training framework. Overall, diffusion-based research in E2E-AD remains limited but holds great potential.
@@ 2.3 Reinforcement Learning for Autonomous Driving
Reinforcement Learning enables an agent to learn an optimal policy through interactions with the environment by maximizing cumulative long-term rewards [41]. Landmark achievements such as AlphaGo [42] and AlphaGo Zero [43] have demonstrated RL’s remarkable capabilities in complex strategy games, while AlphaFold [44] has showcased its power in protein structure prediction. More recently, the successful application of RL in LLMs like DeepSeek-R1 [17] further validates its broad applicability and value. In autonomous driving, several works have begun to explore the integration of RL. Chen et al. [45] propose the ‘world on rails’ assumption, which simplifies RL by decoupling the agent from the environment. GUMP [46] introduces an online training module to improve policy learning efficiency. Lu et al. [47] show that combining imitation learning with simple RL rewards significantly enhances the safety and reliability of learned driving policies. Toromanoff et al. [48] introduce implicit affordances to enable RL to handle urban driving scenarios, including lane keeping, pedestrian and vehicle avoidance, and traffic light interpretation. Zhang et al. [49] train an RL expert to map bird’s-eye view images to low-level continuous actions, providing informative supervision for imitation learning agents. RAD [19] trains an end-to-end driving agent in a photorealistic 3DGS environment using RL. AlphaDrive [18] is among the first to integrate GRPO-based RL with planning and reasoning, substantially improving both performance and training efficiency in E2E-AD.
@@ 3.1 End-to-End Autonomous Driving
Current E2E-AD systems are predominantly based on IL, which aims to learn expert driving policies by mapping raw sensor inputs to ego-vehicle trajectory outputs. It consists of several sequential or unified tasks, including 3D object detection, multi-object tracking, online mapping, motion prediction, and planning. 3D object detection and multi-object tracking tasks provide dynamic semantic features of surrounding agents for planning, while the online mapping module extracts high-confidence lane and road elements as static map features to assist trajectory planning. Among these, planning plays a central role, as it determines the final decision of the ego vehicle.
Multi-Mode Trajectories. The planning module typically predicts a sequence of future waypoints, represented as $\tau=\{(x_{t},y_{t})\}_{t=1}^{T_{f}}$, where $(x_{t},y_{t})$ denotes the ego-vehicle position in its local coordinate frame at time step $t$, and $T_{f}$ is the planning horizon. For datasets such as Bench2Drive[50] and nuScenes [51], trajectories are usually sampled at 2 Hz over a 3-second horizon, resulting in 6 waypoints. To enhance robustness, safety, and adaptability, E2E-AD systems increasingly incorporate diverse driving strategies. A common scheme to enable such diversity is through multi-mode trajectories prediction, where the system predicts $M$ future trajectories, typically $M=6$, denoted as $\tau=\left\{\left(x_{t}^{(m)},y_{t}^{(m)}\right)\,\middle|\,t=1,\dots,T_{f};\;m=1,\dots,M\right\}$. The final trajectory is selected from these candidates based on scoring functions, and the diversity among them (e.g., turning, overtaking, following) is used to infer the ego-vehicle’s potential driving strategies.

### v3

# DIVER: Reinforced Diffusion Breaks Imitation Bottlenecks in End-to-End Autonomous Driving

@@ Abstract
Existing end-to-end autonomous driving (E2E-AD) methods predominantly rely on single expert demonstration through imitation learning, often leading to conservative and homogeneous driving behaviors that struggle to generalize to complex real-world scenarios. In this work, we propose DIVER, a novel E2E-AD framework that combines diffusion-based multi-mode trajectory generation with reinforcement learning to produce diverse, safe, and goal-directed trajectories. First, the model conditions on map elements and surrounding agents to generate multiple reference trajectories from each ground-truth reference trajectory that overcome the inherent limitations of single-mode imitation. Second, we treat the diffusion process as a stochastic policy and employ Group Relative Policy Optimization (GPRO) objectives to guide the diffusion process. By optimizing trajectory-level rewards for both diversity and safety, GRPO directly mitigates mode collapse and enhances collision avoidance, encouraging exploration beyond expert demonstrations and ensuring physically plausible plans. Furthermore, to address the limitations of L2-based open-loop metrics in capturing trajectory diversity, we propose a novel trajectory diversity metric to evaluate the diversity of multi-mode predictions. Extensive experiments on the closed-loop NAVSIM and Bench2Drive benchmarks, as well as the open-loop nuScenes dataset, demonstrate that DIVER significantly improves trajectory diversity, effectively addressing the mode collapse problem inherent in imitation learning. The source code is available at https://github.com/adept-thu/diver.
@@ 1 Introduction
Figure 1: (a) Imitation-based Single-Mode Trajectory Planning [1, 2, 3, 4, 5] predicts deterministic trajectories but lacks action diversity, leading to potential safety risks. (b) Imitation-based Multi-Mode Trajectory Planning [6, 7, 8] fails to address the diversity loss in imitation learning end-to-end autonomous driving, leading to mode collapse. The generated multi-mode trajectories overly depend on a single GT trajectory, ultimately clustering around it. (c) The proposed DIVER framework adopts reinforced diffusion for multi-mode trajectory generation, encouraging the ego-vehicle to produce diverse driving behaviors instead of rigidly following a single expert.
@@
In recent years, autonomous driving has entered a new phase, with significant progress in end-to-end autonomous driving (E2E-AD) [9, 10]. E2E-AD systems integrate traditionally independent tasks, such as 3D object detection, multi-object tracking, online mapping, motion prediction, and planning, into a unified framework that directly learns driving policies from raw sensor inputs [1, 2, 6, 8, 3, 4]. Among these tasks, planning serves as a core module that not only directly governs ego-vehicle decision-making and control but also provides implicit guidance to upstream perception tasks [1].
@@
Currently, imitation learning (IL) remains the dominant paradigm for trajectory planning in E2E-AD systems [9]. IL enables E2E-AD models to mimic human expert driving behaviors through supervised learning, typically by minimizing the discrepancies between predicted and ground-truth (GT) trajectories. The early E2E-AD methods [1, 2, 3, 4] mainly adopted single-mode trajectory generation, as shown in Figure 1(a), where a single deterministic trajectory is regressed toward the expert demonstration. However, such single-mode designs struggle in complex traffic scenarios, as they fail to capture environmental uncertainty and interactions among dynamic agents. To mitigate these limitations, subsequent studies [7, 6, 8, 11] have proposed multi-mode trajectory prediction frameworks, as illustrated in Figure 1(b). For instance, VADv2 [7] employs probabilistic modeling to infer potential driving intentions, generating multiple trajectory proposals and selecting the optimal one via a scoring mechanism. However, these approaches [7, 6, 8, 11] still suffer from the issue of mode collapse, where multi-mode trajectories remain overly concentrated around the GT path, reflecting excessive dependence on expert demonstrations and limited behavioral diversity. This diversity bottleneck stems from the inherent nature of IL, as models are trained to imitate a single expert trajectory in both open-loop and closed-loop settings. As a result, their ability to explore and represent diverse driving behaviors remains fundamentally constrained.
@@
To enhance diversity, several E2E works [12, 13, 8, 14, 15, 16] have recently introduced diffusion models to generate diverse trajectories. Diffusion models [17], as a powerful generative learning framework, have demonstrated strong capabilities in modeling multi-mode behavior distributions for robotic policy learning [18]. Introducing diffusion models into E2E-AD trajectory planning has emerged as a promising research direction. For instance, DiffusionDrive [8] integrates truncated diffusion strategies with efficient cascaded diffusion decoders. VDT-Auto [14] geometrically and contextually parses the environment to condition the diffusion process, while EnDfuser [16] generates a distribution of candidate trajectories from a single perception frame through ensemble diffusion. BridgeDrive [15] further introduces a principled diffusion framework that translates coarse anchors into fine-grained trajectory plans. However, as their optimization processes still adhere to imitation learning (IL) frameworks based on single expert trajectories, the generated multi-mode trajectories ultimately tend to converge toward a single ground truth. This limitation prevents these models from fully overcoming the inherent constraints of imitation learning.
@@
In this work, we propose DIVER, a novel multi-mode E2E-AD framework that leverages reinforcement learning to guide diffusion models for generating diverse, feasible trajectories as shown in Figure 1(c). By leveraging diffusion models, our approach generates multiple plausible future trajectories, capturing a spectrum of driving behaviors while maintaining safety and feasibility. Specifically, we introduce the Policy-Aware Diffusion Generator (PADG), which incorporates map elements and surrounding agents as conditions and generates diverse multi-mode trajectories and multiple reference GTs, guided by a single GT trajectory and trajectory anchors, effectively capturing various driving styles. To ensure both safety and diversity in trajectory planning, we incorporate reinforcement learning to supervise the diffusion process via reward signals, enabling optimization of generated trajectories under diversity and safety constraints. Additionally, to address the misalignment between conventional open-loop metrics (primarily L1 Loss) and the intrinsic requirement for diverse trajectory generation, we introduce a novel Diversity Metric to better evaluate the diversity of multi-mode trajectory predictions. Extensive experiments on closed-loop Bench2Drive and NAVSIM, as well as open-loop nuScenes, demonstrate that our DIVER significantly enhances trajectory diversity and planning robustness, effectively addressing the mode collapse issue in imitation learning.
@@
  • We propose a novel Diversity Metric to evaluate multi-mode trajectory generation, providing a more principled way to assess the diversity and effectiveness of generated trajectories compared to existing metrics.
@@ 2.1 End-to-end Autonomous Driving
End-to-end autonomous driving is a fully differentiable machine learning system that takes raw sensor input data and other metadata as prior information, directly outputting control signals or trajectory planning for vehicles [10, 9]. Early works [19, 20] excluded intermediate tasks, such as perception and motion prediction, which were characterized by limited interpretability and difficulties in optimization. UniAD [1], as a pioneering work, integrates perception tasks and planning tasks via Transformer [21], explicitly supervising 3D object detection [22, 23, 24] , multi-object tracking [25], online mapping [26], motion prediction [27], occupancy prediction [28], and planning [29] tasks within a unified model. Subsequently, VAD [2] further explores the application of vectorized scene representation from MapTR [26] to enhance the efficiency of E2E-AD. DualAD [30] disentangles dynamic agents and static elements, compensating for motion effects and flexibly propagating the belief state over time. A series of E2E-AD works [1, 2, 3, 19, 30, 31], represented by UniAD [1] and VAD [2], have adopted single-mode trajectory for planning performance. However, single-mode trajectory planning methods have issues with action diversity and associated safety risks. To address the above limitations, VADv2 [7] has recently shifted to a multi-mode planning framework by scoring and sampling from an extensive fixed vocabulary of anchor trajectories. Based on the design of multi-mode planning, SparseDrive [6] proposes a hierarchical planning selection strategy to select a rational and safe trajectory as the final planning output. By leveraging historical trajectory momentum, MomAD [32] addresses the issue of inconsistent multi-mode trajectories over time. DiffusionDrive [8] leverages diffusion models to capture multi-mode action distributions in an E2E-AD framework. Existing multi-mode trajectory planning methods [7, 6, 32, 8] do not fundamentally resolve mode collapse in imitation learning. DriveSuprim [33] advances selection-based multi-mode trajectory generation via coarse-to-fine candidate filtering, rotation-based augmentation, and self-distillation, while achieving state-of-the-art safety and trajectory quality. Hydra-MDP [11] uses teacher-student knowledge distillation to incorporate human and rule-based guidance via metric distillation. Overall, our DIVER combines diffusion-based generative modeling with reinforcement learning, thereby bridging generative trajectory modeling and task-aware policy optimization.
@@ 2.2 Diffusion Models for Autonomous Driving
Diffusion models approximate data distributions through iterative denoising and have demonstrated remarkable performance in image generation tasks [34], showing great potential for behavior modeling and trajectory generation in autonomous driving [35]. DiffScene [36] utilizes diffusion models to synthesize safe scenarios for evaluating autonomous driving safety. DiffBEV [37] adopts conditional diffusion to produce fine-grained BEV representations. MotionDiffuser [38] employs diffusion-based representations for multi-agent trajectory prediction. DiffusionDrive [8] models multi-mode trajectory distributions via a truncated diffusion process. VDT-Auto [14] integrates diffusion Transformers with VLMs for action generation. MomAD [32] incorporates noise injection and denoising within the perception module to enhance robustness and generalization. DiffAD [39] formulates E2E-AD as a conditional image generation task by unifying heterogeneous driving objectives into a rasterized representation under a diffusion framework. DiFSD [40] introduces both position-level motion diffusion and trajectory-level planning denoising to model uncertainty, thereby enhancing the stability and convergence of the entire training framework. Overall, diffusion-based research in E2E-AD remains limited but holds great potential.
@@ 2.3 Reinforcement Learning for Autonomous Driving
Reinforcement Learning enables an agent to learn an optimal policy through interactions with the environment by maximizing cumulative long-term rewards [41]. Landmark achievements such as AlphaGo [42] and AlphaGo Zero [43] have demonstrated RL’s remarkable capabilities in complex strategy games, while AlphaFold [44] has showcased its power in protein structure prediction. More recently, the successful application of RL in LLMs like DeepSeek-R1 [45] further validates its broad applicability and value. In autonomous driving, several works have begun to explore the integration of RL. Chen et al. [46] proposed the ‘world on rails’ assumption, which simplifies RL by decoupling the agent from the environment. GUMP [47] introduces an online training module to improve policy learning efficiency. Lu et al. [48] showed that combining imitation learning with simple RL rewards significantly enhances the safety and reliability of learned driving policies. Toromanoff et al. [49] introduced implicit affordances to enable RL to handle urban driving scenarios, including lane keeping, pedestrian and vehicle avoidance, and traffic light interpretation. Zhang et al. [50] trained an RL expert to map bird’s-eye view images to low-level continuous actions, providing informative supervision for imitation learning agents. RAD [51] trains an end-to-end driving agent in a photorealistic 3DGS environment using RL. AlphaDrive [52] is among the first to integrate GRPO-based RL with planning and reasoning, substantially improving both performance and training efficiency in E2E-AD.
@@ 3.1 End-to-End Autonomous Driving
Current E2E-AD systems [1, 2, 6, 8, 3, 4, 5] are predominantly based on IL, which aims to learn expert driving policies by mapping raw sensor inputs to ego-vehicle trajectory outputs. It consists of several sequential or unified tasks, including 3D object detection, multi-object tracking, online mapping, motion prediction, and planning. 3D object detection and multi-object tracking tasks provide dynamic semantic features of surrounding agents for planning, while the online mapping module extracts high-confidence lane and road elements as static map features to assist trajectory planning. Among these, planning plays a central role, as it determines the final decision of the ego vehicle.
Multi-Mode Trajectories. The planning module typically predicts a sequence of future waypoints, represented as $\tau=\{(x_{t},y_{t})\}_{t=1}^{T_{f}}$, where $(x_{t},y_{t})$ denotes the ego-vehicle position in its local coordinate frame at time step $t$, and $T_{f}$ is the planning horizon. For datasets such as Bench2Drive [53] and nuScenes [54], trajectories are usually sampled at 2 Hz over a 3-second horizon, resulting in 6 waypoints. To enhance robustness, safety, and adaptability, E2E-AD systems increasingly incorporate diverse driving strategies. A common scheme to enable such diversity is through multi-mode trajectory prediction, where the system predicts $M$ future trajectories, typically $M=6$, denoted as $\tau=\left\{\left(x_{t}^{(m)},y_{t}^{(m)}\right)\,\middle|\,t=1,\dots,T_{f};\;m=1,\dots,M\right\}$. The final trajectory is selected from these candidates based on scoring functions, and the diversity among them (e.g., turning, overtaking, following) is used to infer the ego-vehicle’s potential driving strategies.